OpenStudy (anonymous):

Evaluate the function rule for the given value. y = 8*3^x , for x = 4 we replace x with 4 8*3^4 = 8*12 = 96

OpenStudy (anonymous):

can u tell me if im right @ospreytriple @mathstudent55

OpenStudy (anonymous):

Not quite. It's important to know the difference between \(3^4\) and \(3 \times 4\). \[3 \times 4 = 12\]while\[3^4 = 3 \times 3 \times 3 \times 3 = 81\]Try again.

OpenStudy (anonymous):

Ohhh okay

OpenStudy (anonymous):

the answer is 648 @ospreytriple

OpenStudy (anonymous):

Correct. Good job.

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

You're welcome.